# USTC-ICTS/PCFT-25-07

### v1

@@ Abstract
General relativity (GR) provided us a picture that gravitation is identical to curved spacetime, and is formulated by Riemannian geometry. With this, gravity is manifested by the curvature which is constructed from the Levi-Civita connection (or Christoffel symbol) and in turn from the metric. However, Riemannian geometry is not the unique approach to gravity theories. Sample theories basing on non-Riemannian geometries include the Einstein-Cartan theory, the metric-affine theory, the teleparallel gravity, and so on. Usually in these theories, the connection is not limited to be the Levi-Civita type, and has no a priori dependence on the metric. For the teleparallel gravity Tele ; Bahamonde:2021gfp ; Krssak:2018ywd , the curvature obtained from the connection vanishes and gravity is manifested by other geometric quantities: torsion in the metric teleparallel gravity (MTG), or non-metricity in the symmetric teleparallel gravity (STG) Nester:1998mp . Within both frameworks one can build models equivalent to GR. This means treating the same thing in different pictures.
@@
This paper is organized as follows: First in we will briefly introduce the teleparallel gravity in Section II, then discuss the ways of formulating the spacetime perturbations with non-metricity and torsion respectively in Section III and IV, and conclude in Section V.
@@ II Teleparallel gravity
where $T_{\mu}=T^{\nu}_{~{}\nu\mu}$ is the torsion vector and the defined $\mathbb{T}$ is a torsion scalar and is in quadratic form of the torsion tensor. With these, one can construct a model equivalent to general relativity within the framework of MTG. This model, dubbed TGR, has the action
@@
it is the same as $S_{\rm GR}$ after integrating out the total derivative term in Eq. (13). Other MTG models are considered as extensions to TGR, such as the $f(\mathbb{T})$ model Yang:2010ji ; Cai:2015emx ; Krssak:2015oua ; Bahamonde:2022ohm and the Nieh-Yan modified Teleparallel Gravity (NYTG) model Li:2020xjt ; Li:2021wij ; Li:2023fto ; Zhang:2024vfw , and so on. The MTG model gives a picture that gravity is manifested by torsion in stead of curvature. Please note that in MTG, the affine connection $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ or the torsion tensor are not fundamental variables. The teleparallelism constraint determined that $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ can neither have a general form nor be varied freely when using the variation principle. Furthermore, torsion is required to be existent, so that one cannot simply set all the components of $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ to zero to fit the teleparallelism constraint. To find the true building blocks of the MTG models, it is convenient to make use of the tetrad formulation. With this language the metric is built from the the tetrad (or veilbein) $e^{a}_{\mu}$ through the relation $g_{\mu\nu}=e^{a}_{\mu}e^{b}_{\nu}\eta_{ab}$, here $\eta_{ab}={\rm diag}(-1,+1,+1,+1)$ is the Minkowski metric and the Latin letters $a,b,...$ are the Lorentz indices, used to denote the tensor components at the local flat space. The affine connection $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ is constructed by the tetrad and the spin connection $\hat{\omega}^{a}_{~{}b\mu}$ as $\hat{\Gamma}^{\rho}_{~{}\mu\nu}=\theta^{\rho}_{a}(\partial_{\mu}e^{a}_{\nu}+\hat{\omega}^{a}_{~{}b\mu}e^{b}_{\nu})$, here $\theta^{\rho}_{a}$ is the inverse of the tetrad: $\theta^{\rho}_{a}e^{a}_{\sigma}=\delta^{\rho}_{\sigma}$ and $\theta^{\rho}_{a}e^{b}_{\rho}=\delta^{b}_{a}$. Then with the definition $T^{\rho}_{~{}\mu\nu}=\hat{\Gamma}^{\rho}_{~{}\mu\nu}-\hat{\Gamma}^{\rho}_{~{}\nu\mu}$, it is straightforwardly to obtain the expression of the torsion tensor
@@
where $Q^{\mu}=Q^{\mu\nu}_{~{}~{}~{}\nu}~{},~{}\bar{Q}^{\mu}=Q^{\nu~{}\mu}_{~{}\nu}$ are two non-metricity vectors, and the defined $\mathbb{Q}$ is a non-metricity scalar which is in quadratic form of the non-metricity tensor. Similarly, one can construct a model equivalent to general relativity within the framework of STG. This model, dubbed QGR, has the action
@@
From now on, we address to the question of how to formulate gravitations from the background and the perturbation with separate pictures in the perturbation theory. We first consider this problem within the metric formulation. With this language, we have a metric $g_{\mu\nu}$ for the physical spacetime and its Levi-Civita connection $\Gamma^{\rho}_{~{}\mu\nu}$ defined in Eq. (1). At the same time we have a metric $\bar{g}_{\mu\nu}$ for the background spacetime and its corresponding Levi-Civita connection $\bar{\Gamma}^{\rho}_{~{}\mu\nu}=\frac{1}{2}\bar{g}^{\rho\sigma}(\partial_{\mu}\bar{g}_{\sigma\nu}+\partial_{\nu}\bar{g}_{\mu\sigma}-\partial_{\sigma}\bar{g}_{\mu\nu})$. Both the physical and background spacetimes are Riemannian geometric, and manifest respective gravitational interactions through respective curvatures, $R(\Gamma)$ and $\bar{R}(\bar{\Gamma})$. The spacetime perturbation arise from the difference between these two metrics, this leads to the mismatch between $g_{\mu\nu}$ and $\bar{\Gamma}^{\rho}_{~{}\mu\nu}$, or between $\bar{g}_{\mu\nu}$ and $\Gamma^{\rho}_{~{}\mu\nu}$.
@@
where $\epsilon$ is a small matrix and its elements $\epsilon^{\mu}_{~{}\nu}$ are considered as the basic perturbation variables. Hence one can obtain that
@@ V Conclusions
Gravity is identical to curved spacetime and can be manifested by curvature, torsion or non-metricity. Armed with these multiple options, we in this paper revisited the problem of separating the physical spacetime into background and perturbation in perturbation theory, and considered the possibilities of formulating the gravitation of background and that of perturbation in separated ways. We showed that the perturbation to the curvature of a Riemannian spacetime can be represented in terms of non-metricity (in the metric formulation) or torsion (in the tetrad formulation), but the background is still Riemannian geometric. With these separate treatments, we got teleparallel like actions for the spacetime perturbation around a Riemannian background.

### v2

@@ Abstract
General relativity (GR) provided us a picture that gravitation is identical to curved spacetime, and is formulated by Riemannian geometry. With this, gravity is manifested by the curvature which is constructed from the Levi-Civita connection (or Christoffel symbol) and in turn from the metric. However, Riemannian geometry is not the unique approach to gravity theories. Sample theories based on non-Riemannian geometries include the Einstein-Cartan theory, the metric-affine theory, the teleparallel gravity, and so on. Usually in these theories, the connection is not limited to be the Levi-Civita type, and has no a priori dependence on the metric. For the teleparallel gravity Tele ; Bahamonde:2021gfp ; Krssak:2018ywd , the curvature obtained from the connection vanishes and gravity is manifested by other geometric quantities: torsion in the metric teleparallel gravity (MTG), or non-metricity in the symmetric teleparallel gravity (STG) Nester:1998mp . Within both frameworks one can build models equivalent to GR. This means treating the same thing in different pictures.
@@
This paper is organized as follows: First we will briefly introduce the teleparallel gravity in Section II, then discuss the ways of formulating the spacetime perturbations with non-metricity and torsion respectively in Section III and IV, and conclude in Section V.
@@ II Teleparallel gravity
where $T_{\mu}=T^{\nu}_{~{}\nu\mu}$ is the torsion vector and the defined $\mathbb{T}$ is a torsion scalar and is in quadratic form of the torsion tensor. With these, one can construct a model equivalent to general relativity within the framework of MTG. Such a model depends heavily on torsion tensor (denoted by $T$) and is equivalent to general relativity, we may call it TGR model. It has the action
@@
it is the same as $S_{\rm GR}$ after integrating out the total derivative term in Eq. (13). Other MTG models are considered as extensions to TGR, such as the $f(\mathbb{T})$ model Yang:2010ji ; Cai:2015emx ; Krssak:2015oua ; Bahamonde:2022ohm and the Nieh-Yan modified Teleparallel Gravity (NYTG) model Li:2020xjt ; Li:2021wij ; Li:2023fto ; Zhang:2024vfw , and so on. The MTG model gives a picture that gravity is manifested by torsion in stead of curvature. Please note that in MTG, the affine connection $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ or the torsion tensor are not fundamental variables. The teleparallelism constraint determines that $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ can neither have a general form nor be varied freely when using the variation principle. Furthermore, torsion is required to be existent, so that one cannot simply set all the components of $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ to zero to fit the teleparallelism constraint. To find the true building blocks of the MTG models, it is convenient to make use of the tetrad formulation. With this language the metric is built from the the tetrad (or veilbein) $e^{a}_{\mu}$ through the relation $g_{\mu\nu}=e^{a}_{\mu}e^{b}_{\nu}\eta_{ab}$, here $\eta_{ab}={\rm diag}(-1,+1,+1,+1)$ is the Minkowski metric and the Latin letters $a,b,...$ are the Lorentz indices, used to denote the tensor components at the local flat space. The affine connection $\hat{\Gamma}^{\rho}_{~{}\mu\nu}$ is constructed by the tetrad and the spin connection $\hat{\omega}^{a}_{~{}b\mu}$ as $\hat{\Gamma}^{\rho}_{~{}\mu\nu}=\theta^{\rho}_{a}(\partial_{\mu}e^{a}_{\nu}+\hat{\omega}^{a}_{~{}b\mu}e^{b}_{\nu})$, here $\theta^{\rho}_{a}$ is the inverse of the tetrad: $\theta^{\rho}_{a}e^{a}_{\sigma}=\delta^{\rho}_{\sigma}$ and $\theta^{\rho}_{a}e^{b}_{\rho}=\delta^{b}_{a}$. Then with the definition $T^{\rho}_{~{}\mu\nu}=\hat{\Gamma}^{\rho}_{~{}\mu\nu}-\hat{\Gamma}^{\rho}_{~{}\nu\mu}$, it is straightforwardly to obtain the expression of the torsion tensor
@@
where $Q^{\mu}=Q^{\mu\nu}_{~{}~{}~{}\nu}~{},~{}\bar{Q}^{\mu}=Q^{\nu~{}\mu}_{~{}\nu}$ are two non-metricity vectors, and the defined $\mathbb{Q}$ is a non-metricity scalar which is in quadratic form of the non-metricity tensor. Similarly, one can construct a model equivalent to general relativity within the framework of STG. Such a model depends heavily on non-metricity tensor (denoted by $Q$) and is equivalent to general relativity, we may call it QGR model. It has the action
@@
From now on, we address to the question of how to formulate gravitations from the background and the perturbation with separate pictures in the perturbation theory. We first consider this problem within the metric formulation. With this language, we have a metric $g_{\mu\nu}$ for the physical spacetime and its Levi-Civita connection $\Gamma^{\rho}_{~{}\mu\nu}$ defined in Eq. (1). At the same time we have a metric $\bar{g}_{\mu\nu}$ for the background spacetime and its corresponding Levi-Civita connection $\bar{\Gamma}^{\rho}_{~{}\mu\nu}=\frac{1}{2}\bar{g}^{\rho\sigma}(\partial_{\mu}\bar{g}_{\sigma\nu}+\partial_{\nu}\bar{g}_{\mu\sigma}-\partial_{\sigma}\bar{g}_{\mu\nu})$. Both the physical and background spacetimes are of Riemannian geometries, and manifest respective gravitational interactions through respective curvatures, $R(\Gamma)$ and $\bar{R}(\bar{\Gamma})$. The spacetime perturbation arise from the difference between these two metrics, this leads to the mismatch between $g_{\mu\nu}$ and $\bar{\Gamma}^{\rho}_{~{}\mu\nu}$, or between $\bar{g}_{\mu\nu}$ and $\Gamma^{\rho}_{~{}\mu\nu}$.
@@
where $\epsilon$ is a small matrix and its elements $\epsilon^{\mu}_{~{}\nu}$ are considered as the basic perturbation variables. This parameterization of the perturbation is not conventional, but for our purposes in this paper it has some advantages relative to the conventional parameterization of the metric perturbation, such as $g_{\mu\nu}=\bar{g}_{\mu\nu}+h_{\mu\nu}$ or the exponential form $g_{\mu\nu}=\bar{g}_{\mu\rho}\left(e^{h}\right)^{\rho}_{~{}\nu}$. With the parameterization of Eq. (22), $e^{\epsilon}$ has the meaning of transfer matrix for the map between the physical spacetime and the background. It is the same with the parameterization of the tetrad perturbation, which will be discussed in the next section. In addition, this parameterization automatically guaranteed the symmetric property of $g_{\mu\nu}$ as long as the background metric $\bar{g}_{\rho\sigma}$ is symmetric. As a comparison, in the conventional method, one should presuppose the symmetry of $h_{\mu\nu}$. But with the parameterization of Eq. (22), the linear perturbation to the metric is $\epsilon_{\mu\nu}+\epsilon_{\nu\mu}$ with $\epsilon_{\mu\nu}\equiv\bar{g}_{\mu\rho}\epsilon^{\rho}_{~{}\nu}$, which corresponds to $h_{\mu\nu}$ in the conventional parameterization. With Eq. (22), one can obtain that
@@ V Conclusions
Gravity is identical to curved spacetime and can be manifested by curvature, torsion or non-metricity. Armed with these multiple options, we in this paper revisited the problem of separating the physical spacetime into background and perturbation in perturbation theory, and considered the possibilities of formulating the gravitation of background and that of perturbation in separated ways. We showed that the perturbation to the curvature of a Riemannian spacetime can be represented in terms of non-metricity (in the metric formulation) or torsion (in the tetrad formulation), but the background is still of Riemannian geometry. With these separate treatments, we got teleparallel like actions for the spacetime perturbation around a Riemannian background.
Torsion and non-metricity have applications in other branches of physics. In the solid physics, topological defects caused by plastic deformations to the ideal crystals can be also formulated in the differential geometry. In this language, a kind of linear defects, called dislocations, is described by torsion (see the section 3.9 of Ref. Bahamonde:2021gfp ), and the point-like defects are described by non-metricity (see Ref. pointlike for an example). As an analogy, in this paper we want to provide a preliminary image in which the spacetime perturbations are considered as point-like topological defects or dislocations randomly distributed over the background spacetime.
This formalism can be extended to some more general theories, such as the scalar-tensor theories in which the Lagrangian density is a linear function of the Ricci scalar, or the $f(R)$ theory which is equivalent to the former case after Legendre transformation and field redefinition. These theories were originally formulated in the so-called Jordan frame. However, through conformal transformation (or Weyl rescaling) they can be transformed into the Einstein frame in which the action of gravity has the form of Einstein-Hilbert, taking the price of introducing non-minimal couplings to the matter sector. Then the formulation presented here can be applied to these theories in strait forward ways.